# Improve Value Estimation of Q Function and Reshape Reward with Monte Carlo Tree Search

### v1

@@ V Conclusion
We introduced the complexity of imperfect information games and discussed their research value. We selected Uno, a imperfect information game, as the basis for our research. We represented Uno as a reinforcement learning problem and presented a novel algorithm, Double Deep Q-Learning with Monte Carlo Tree Search (DDQN with MCTS), to address challenges encountered in prior work with imperfect information games, such as reward sparsity and Q-value overestimation. Additionally, we developed a graphical user interface (GUI) to allow human players to compete against our agents, where DDQN with MCTS outperformed the average human player in terms of win rates on one-to-one competition. DDQN with MCTS also demonstrated superior performance compared to three traditional methods—Double Deep Q-Learning, Deep Monte Carlo, and Neural Fictitious Self-Play—achieving higher total average rewards with fewer training steps and higher win rates during testing.

### v2

@@ V Conclusion
We introduced the complexity of imperfect information games and discussed their research value. We selected Uno, an imperfect information game, as the basis for our research. We represented Uno as a reinforcement learning problem and presented a novel algorithm, Double Deep Q-Learning with Monte Carlo Tree Search (DDQN with MCTS), to address challenges encountered in prior work with imperfect information games, such as reward sparsity and Q value overestimation. Additionally, we developed a graphical user interface (GUI) to allow human players to compete against our agents, where DDQN with MCTS outperformed the average human player in terms of win rates on one-to-one competition. DDQN with MCTS also demonstrated superior performance compared to three traditional methods—Double Deep Q-Learning, Deep Monte Carlo, and Neural Fictitious Self-Play—achieving higher total average rewards with fewer training steps and higher win rates during testing.
@@
Acknowledgments
We would like to thank Amitabh Trehan for his advice in the development of this research.